# IA-TIGRIS: An Incremental and Adaptive Sampling-Based Planner for Online Informative Path Planning

### v1

@@ 12             $V_{closed}\xleftarrow{+}\{n_{new}\}$
IV-B Adaptive Replanning
@@
For environments where the world is static and measurements are deterministic, global plans could be static and created once before flight. Having a single, static plan would also work well in situations when onboard computation is restricted, there is high confidence the global path is optimal, the executed path will be perfectly followed, and the predicted information gain will be close to the actual information gain. However, there are many cases where the plan would need to be adjusted due to disturbances and modified due to new information. When disturbances cause deviations from an expected observation, the path can be replanned during path execution to re-observe locations and adjust the path according to the current information map.
@@
The adaptability of IA-TIGRIS is due to its ability to continuously replan the global trajectory online. The updated paths take into account the current world belief updated by all observations. This brings robustness, as executed paths may not follow the planned path due to disturbances and lead to differences in the planned versus expected observations.
@@ IV-C Incremental Planning
When replanning online, discarding the previous planning graph wastes valuable information and forces the planner to start from scratch. By using an incremental approach, the planner can reuse the previous planning efforts, improving efficiency and enabling the next global plan to refine the earlier one. This often results in higher-quality paths. Even when the budget and belief map remain unchanged between replans, many scenarios—such as large budgets or small areas relative to the budget—can benefit from incremental refinement. In challenging planning situations, this approach allows the planner to iteratively expand the search tree, continuing to explore and evaluate potential paths.
To enable incremental planning, our method recycles as much of the previously-built graph as possible, rather than create a new graph whenever a plan is requested. Even when the belief map and budget have changed, recycling the previous graph offers significant benefits. The previous graph has already been validated for collisions and adheres to the vehicle’s motion model, allowing the planner to build upon prior work and save computational resources while maintaining accuracy.
@@
To reuse the previous tree, Algorithm 1 first checks if the vertex set $V$ is empty. If it is not empty, then rather than initializing the graph, it calls the function UpdateGraph. This function first finds where the starting point for the next plan lies within the previous plan. When a matching node is found in the previous plan, all nodes and edges before that node are pruned away. The remaining portions of the graph are checked for the budget constraint and the information gain up to each node are updated based on the current information map. This is done by efficiently traversing the tree recursively from the starting node working down to all leaf nodes. The planner then uses the updated graph to plan as normal. A high-level overview of this process is shown in Fig. 4.
IV-D Belief Space Representation and Node Embedding
In informative path planning, the information within a space can be represented using various models, such as Gaussian mixture models, Gaussian process regressions, grids, or voxels. This belief space can encode different types of information. For example, it might represent the probability of an object of interest being located in specific regions of the space. The choice of belief space formulation is important, as it directly influences the planning process by guiding search efforts and behaviors. An effective representation for a specific system, application, and planner balances computational complexity, memory requirements, and the ability to perform operations on the belief model.
In this work, we adopt a grid-based representation of the belief space. Specifically, the search space is discretized into a grid of cells, where each cell $i$ is associated with a binary random variable $X_{i}$ indicating the presence ($X_{i}=1$) or absence ($X_{i}=0$) of an object of interest. The probability $P(X_{i})$ represents our belief about an object being located in cell $i$. This grid-based representation is computationally straightforward to update with new or predicted observations using a Bayesian framework and an appropriate sensor model.
IV-D1 Belief Space Node Embedding
The nodes in the planning tree maintained by the IA-TIGRIS algorithm store the necessary information to represent the evolving belief space. The information gain up to a given node is computed based on the change in entropy within the grid cells observed by the agent’s sensor footprint during traversal from the root node to the current node.
A straight-forward approach to calculating the information gain for a given node consists of traversing the tree from the root all the way to the given node. It is necessary to evaluate the entire trajectory because the information gain can be not modular and previous observations affect the information gain of future observations. Calculating the additional information gain contributed by a new node and edge requires knowledge of the belief state of the search map at the previous node. However, this approach of calculating information gain starting at the root node is computationally inefficient and slow. Let $n$ be the number of cells in the search map, $m$ be the number of nodes in the tree, $k$ be the number of nodes in a trajectory, and $j$ be the upper bound of cells updated by a node in the tree. The temporal complexity of this method is $\mathcal{O}(kj)$ and the memory complexity is $\mathcal{O}(n)$.
One approach to speeding up computation is to store the entire search map within each node. In this approach, the information gain for a new node is simply the sum of the parent node’s information gain and the contribution of the new node. This method results in a temporal complexity of $\mathcal{O}(j)$ for calculating the information gain of a new node and a memory complexity of $\mathcal{O}(nm)$. While feasible for small maps, this method becomes impractical for large-scale maps relevant to the domains and applications considered in this work due to its high memory requirements.
To efficiently compute the information gain for a new node, we introduce a node embedding strategy that eliminates the need for a full traversal of the planning tree from the root to the current node. Instead of storing the complete belief state in each node, we use a sparse representation that records only belief updates. A hash map efficiently tracks modifications to belief probabilities, reducing memory complexity while preserving the ability to compute information gain incrementally. This allows for efficient calculations of the information gain for a new node while maintaining computational efficiency without sacrificing accuracy.
The temporal complexity of calculating the information gain for a new node is now $\mathcal{O}(j)$, and the overall memory complexity of this belief space representation is only $\mathcal{O}(jm+n)$. As long as $j\ll n$, this approach is much more memory-efficient than the naive approach of storing the complete belief state in each node. Each node in our tree now contains the cumulative information gain along the path, current state of the agent, path cost, and edge path.
While implementing this embedded method, we observed the significant impact on computation time from selecting efficient keys for the hashmap. Using strings as keys resulted in substantial overhead due to frequent string-to-integer conversions. Even transitioning from a two-integer key to a single-integer key yielded a notable improvement in performance and is the approach we use in our implementation through a simple index-to-key conversion. Given that cell value lookup operation occurs frequently, optimizing its efficiency was highly impactful for accelerating the expansion rate of the search tree.
@@ IV-D1 Belief Space Node Embedding
Figure 4: An illustration of the incremental planning and refinement process in IA-TIGRIS. (a) The global planner first generates long-horizon paths to maximize predicted information gain. (b) At the start of each new planning cycle, infeasible portions of the tree, based on the executed trajectory and the current plan, are pruned. The remaining tree is then updated by recomputing information gain and cost based on the latest belief state and available budget, which is significantly more efficient than rebuilding the tree from scratch. (c) The tree is further expanded and refined for the duration of the planning cycle, after which the process repeats.
@@
By incorporating this belief space node embedding strategy, our planning framework efficiently maintains an accurate representation of the evolving belief space within the search tree. This enables the planner to greatly reduce computational overhead and create higher quality paths.
@@ IV-E Sensor Modeling
The updates to our belief map are dependent on the sensor model, which captures the performance characteristics of the perception system employed. There are many environmental factors that affect the performance of a sensor, such as visibility and range, which can be incorporated into this model.
For our approach, we formulate the sensor model as a piecewise function and is implemented as a lookup table. For a given range $r$ and binary measurement $Z$, this table maps the range of the object from the sensor to the corresponding true positive rate (TPR) $P\left(Z|X,r\right)$ and true negative rate (TNR) $P\left(\overline{Z}|\overline{X},r\right)$ of an object detection model. A typical range-based detection model will have better performance at close ranges and then tapers off as the distance increases.
@@ IV-F Information Reward Models
The information reward function $I(\cdot)$ used in our framework is not constrained to being modular and could be time-varying or even submodular, depending on the specific requirements of the application. In this work, we have implemented an objective of reducing the overall entropy of the belief space. As defined in Section IV-D, our belief space represents the probability of the presence of an object of interest in a cell. The reward for one cell $I(X_{i})$ is calculated by finding the difference between the initial entropy and the final entropy after updating the belief with the measurements and sensor model defined in Section IV-E.
@@ IV-F2 Priority-dependent Rewards
In the case where certain areas of an environment or spatial information are more important than others, we introduced the ability to weight the information reward by a priority scalar, $p_{i}$, representing the priority of cell $i$. This approach provides finer control over the UAV behavior by allowing the entropy of different cells to be weighted unequally. Assigning a higher priority to one prior over another results in paths that favor observations of the higher-weighted prior. For example, this is useful when one type of object is more important than another, and the prior spatial distributions are concentrated in distinct areas of the search space (as demonstrated in Section V-C3).
@@ IV-F3 Time-dependent Rewards
In our use cases, we implement a simple reward decay function
@@
where $\Gamma(t)\in[\gamma,1]$ and $\gamma$ is the floor value of the decay function. $\beta$ is the decay rate and $\gamma_{t}=(\gamma-1)/\beta$ is the time where the function output becomes the floor value $\gamma$. This gives a linear decay until the floor value.
@@ V Simulation Evaluation & Results
Figure 7: IA-TIGRIS continuously replanning and refining the global path over the duration of the simulation. These snapshots show how the path is refined online to maximize information gain. The search tree is also visualized and shows how the plans focus on the areas of high information gain and take into account the remaining budget.
@@ V-A Simulation Framework
We use a simplified simulator, shown in Fig. 6, to enable rapid development and validation of our approach. The simulation employs a constant-velocity motion model for the fixed-wing UAV. To account for the effects of banking on observations, we compute the orientation change between consecutive waypoints. If this change exceeds a threshold, the UAV is considered to be banking, and the belief space is not updated during this period. This reflects the fact that the camera is primarily pointed toward the sky during banking and is also affected by increased motion blur.
@@
The UAV commanded speed is $25$ m/s, with a planning time of $10$ seconds and a the flight altitude of $50$ m. The turning radius is $100$ m, and the onboard camera is pitched $30$ degrees from horizontal with a $36.9$-degree field of view. The belief space grid cells are $30$ m, and the total budget for flight distance is $15,000$ m. To accelerate testing, the simulation runs at twice the real-time speed, with results subsequently rescaled to match real-time conditions.

### v2

@@ 12        $V_{closed}\xleftarrow{+}\{n_{new}\}$
At the beginning of planning in Algorithm 1, the set of vertices $V$ is empty. The initial information $I_{s}$ at the start pose $x_{start}$ is calculated by $I(\cdot)$, and the initial cost $C$ is set to 0 (Line 2). The tuple $n_{start}$ is initialized and added to the set of vertices $V$, while the set of edges $E$ and the closed set of vertices $V_{closed}$ are initialized as empty (Lines 3-4). The graph G is initialized as $G\leftarrow(V,E)$.
The algorithm then loops until the computation time exceeds the allowed planning time $T$ (Line 9). Each iteration draws a sample $x_{sample}$ from the state space $\mathcal{X}$ using InformedSample, which is a weighted sampler based on the reward for viewing a state in the belief map. This is done by first conducting the weighted sample in the belief map, then uniformly sampling in the subset of the configuration space that views the sampled belief map state while respecting desired observation constraints (see [21] for related details). Informed sampling focuses paths on areas of higher reward rather than uniformly sampling in the search space.
The algorithm identifies the nearest node $n_{nearest}$ in the graph $G$ to $x_{sample}$ (Line 11) and attempts to steer from $n_{nearest}$ toward $x_{sample}$ within distance $\Delta$ while checking for collisions, resulting in a feasible pose $x_{feas}$ and an edge $e_{feas}$ (Line 12). Unlike [21], the new pose and edge are immediately evaluated for addition to the graph $G$ using AddPoseToGraph (Line 13), allowing the neighbor selection radius $R$ to be less than the extend distance $\Delta$.
The algorithm then searches for nearby nodes $N_{near}$ within a radius $R$ of $x_{feas}$ that are not in the closed set $V_{closed}$ (Line 14). For each neighbor $n_{near}$ in $N_{near}$, if $n_{near}$ is distinct from $x_{feas}$ and $n_{nearest}$, the algorithm attempts to steer from $n_{near}$ to $x_{feas}$ (Lines 15-17). If the new pose $x_{new}$ is distinct from $n_{near}$, the pose and edge evaluated to be added to the graph $G$ using the AddPoseToGraph function (Lines 18-19). This neighborhood expansion structure is different than [21] in that it now allows for neighboring nodes to extend toward $x_{feas}$ even if it did not add an edge connecting to $n_{nearest}$ due to a collision or budget constraint.
After the computation time is exhausted, the algorithm extracts the best path $T$ from the graph $G$ using the BestPath function (Line 24). This function simply returns the path with the highest information reward in the graph.
The AddPoseToGraph function, outlined in Algorithm 2, handles the addition of a new pose $x_{new}$ and edge $e_{new}$ to the graph $G$. The information $I_{new}$ for the new pose is calculated, and the cumulative cost $C_{new}$ is computed by adding the cost of the edge $e_{new}$ to the cost of $n_{parent}$ (Lines 1-2). A new node $n_{new}$ is created as a tuple containing the new pose, information, and cumulative cost (Line 3). The node $n_{new}$ is then checked by the Prune function (Line 4). The pruning function checks nearby nodes within a radius to see if they contain a better solution. Because most of the nodes in the tree are not full paths, the upper bound on their rewards are compared. For more aggressive pruning, heuristics can be used for this comparison. In our case we use the heuristic of checking if a nearby node has a lower cost and higher current reward, but this heuristic can have implications on finding the optimal solution. If $n_{new}$ passes the pruning check, the edge $e_{new}$ and node $n_{new}$ are added to the sets $E$ and $V$, respectively, and graph $G$ is updated (Lines 4-7). If the cumulative cost $C_{new}$ equals the budget $B$, the new node $n_{new}$ is added to the closed set $V_{closed}$ (Lines 8-9).
IV-B Adaptive Planning
@@
For environments where the world is static and measurements are deterministic, global plans could be static and created once before flight. Having a single, static plan would also work well in situations when onboard computation is restricted, there is high confidence the global path is optimal, the executed path will be perfectly followed, and the predicted information gain will be close to the actual information gain. However, there are many cases where the plan would need to be adjusted due to disturbances and modified due to new information. When disturbances cause deviations from an expected observation, the path can be replanned during path execution to re-observe locations and adjust the path according to the current belief map.
@@
The high adaptability of IA-TIGRIS is due to its ability to continuously replan the global trajectory online. The updated paths take into account the current world belief updated by all observations. This brings robustness, as executed paths may not follow the planned path due to disturbances and lead to differences in the planned versus expected observations.
1
2
Input : $\mathbf{x}_{start}$: New starting pose
$B$: Budget
3
[…]
5
6if $n_{match}$ exists then
7    PruneBefore($n_{match}$)
8    UpdateSubtree($n_{match}$, $B$)
9    $n_{start}\leftarrow n_{match}$
10 end if
11else
12    $I_{s}\leftarrow I(\mathbf{x}_{start})$; $C\leftarrow 0$
13    $n_{start}\leftarrow\{\mathbf{x}_{start},I_{s},C\}$
15
16 end if
Algorithm 3 UpdateGraph
@@ IV-C Incremental Planning
Rebuilding the planning graph from scratch for every replan discards valuable information. In contrast, an incremental approach allows the planner to reuse previous planning efforts. Many scenarios, such as large budgets or small areas relative to the budget, can benefit from incremental refinement. This not only improves the planner efficiency but also allows each new global plan to build on and refine earlier ones. In challenging planning situations, this approach allows the planner to iteratively expand the search tree, continuing to explore and evaluate potential paths over time. Even when the budget and belief map remain unchanged between replans, the previous graph remains valuable as it has already been validated for collisions and adheres to the vehicle’s motion model, allowing the planner to build upon prior work and save computational resources while maintaining accuracy.
@@
To reuse the previous tree for incremental planning, Algorithm 1 first checks if the vertex set $V$ is empty. If it is not, rather than initializing the graph, it calls the function UpdateGraph, outlined in Algorithm 3. This function first finds where the new starting point for the next plan matches a node in the existing previous trajectory (Line 1). When a matching node is found, all nodes and edges before that node are pruned away (Lines 2-3). The remaining portions of the graph are updated by checking the budget constraint and recomputing the information gain up to each node based on the current belief map. This is done efficiently by recursively traversing the tree using the helper function UpdateSubtree (Line 4), starting from the matching node and working down to all leaf nodes. The planner then uses the updated graph to continue planning as normal. If no matching node is found, the planner simply initializes a new graph from the new start pose (Lines 7-10). A high-level overview of this process is shown in Fig. 4.
@@
Figure 4: An illustration of the incremental planning and refinement process in IA-TIGRIS. (a) The global planner first generates long-horizon paths to maximize predicted information gain. The blue ellipses correspond to areas of information where areas with a larger alpha correspond to higher information. (b) At the start of each new planning cycle, infeasible portions of the tree, based on the executed trajectory and the current plan, are pruned. The remaining tree is then updated by recomputing information gain and cost based on the latest belief state and available budget, which is significantly more efficient than rebuilding the tree from scratch. (c) The tree is further expanded and refined for the duration of the planning cycle, after which the process repeats.
IV-D Belief Map Representation and Node Embedding
In informative path planning, the information within a space can be represented using various models, such as Gaussian mixture models, Gaussian process regressions, grids, or voxels. This belief map can encode different types of information. For example, it might represent the probability of an object of interest being located in specific regions of the space. The choice of belief map formulation is important, as it directly influences the planning process by guiding search efforts and behaviors. An effective representation for a specific system, application, and planner balances computational complexity, memory requirements, and the ability to perform operations on the belief model.
In this work, we adopt a grid-based representation of the belief map. Specifically, the search space is discretized into a grid of cells, where each cell $i$ is associated with a binary random variable $X_{i}$ indicating the presence ($X_{i}=1$) or absence ($X_{i}=0$) of an object of interest. The probability $P(X_{i})$ represents our belief about an object being located in cell $i$. This grid-based representation is computationally straightforward to update with new or predicted observations using a Bayesian framework and an appropriate sensor model.
While we adopt a binary occupancy grid here to represent the presence of discrete objects, other belief models can capture spatial correlations and are better suited for continuous-valued fields like temperature or gas concentration. These richer models, however, typically increase computational cost. In our target applications, where objects of interest are assumed to be independent and onboard real-time planning is critical, the grid-based model is a practical choice. Importantly, our planner itself is agnostic to the belief representation and could be adapted to different models as needed.
IV-D1 Belief Map Node Embedding
The nodes in the planning tree maintained by the IA-TIGRIS algorithm store the necessary information to represent the evolving belief map. The information gain up to a given node is computed based on the change in entropy within the grid cells observed by the agent’s sensor footprint during traversal from the root node to the current node.
A straight-forward approach to calculating the information gain for a given node consists of traversing the tree from the root all the way to the given node. It is necessary to evaluate the entire trajectory because the information gain can be not modular and previous observations affect the information gain of future observations. Calculating the additional information gain contributed by a new node and edge requires knowledge state of the belief map at the previous node. However, this approach of calculating information gain starting at the root node is computationally inefficient and slow. Let $n$ denote the number of grid cells in the belief map, $m$ the total number of nodes in the planning tree, $k$ the number of nodes in the trajectory from the root to the current node, and $j$ an upper bound on the number of cells affected by a single sensor footprint. Computing the information gain for the current node entails stepping through each of the $k$ nodes along the trajectory and, at every step, updating at most $j$ cells, giving a running time of $\mathcal{O}(kj)$. Because the belief map is maintained once and updated in place during this traversal, the memory footprint remains $\mathcal{O}(n)$.
One approach to speeding up computation is to store the entire belief map within each node. In this approach, the information gain for a new node is simply the sum of the parent node’s information gain and the contribution of the new node. With this design, evaluating a newly added node involves touching only the $j$ cells observed at that node, so the per-update cost drops to $\mathcal{O}(j)$. The trade-off is a higher space requirement: each of the $m$ nodes now holds $n$ cell probabilities, for a total memory complexity of $\mathcal{O}(mn)$. While feasible for small maps, this method becomes impractical for large-scale maps relevant to the domains and applications considered in this work due to its high memory requirements.
To efficiently compute the information gain for a new node, we introduce a node embedding strategy that eliminates the need for a full traversal of the planning tree from the root to the current node, by embedding only the delta of the belief state at each node. Concretely, every node stores the subset of grid map cells—no more than $j$ of them—that its sensor footprint updates. This subset is stored as a hash map, keyed by cell index, maintaining the most recent probability of the visited cell; unvisited cells fall back to the baseline grid that is held once for the whole problem. When a new node is appended, the planner consults the hash map for at most $j$ entries, updates those values, and aggregates the resulting change in entropy into the node’s cumulative information-gain value, so the per-node running time is $\mathcal{O}(j)$.
Because each node retains at most $j$ key–value pairs, the additional storage it contributes is $\mathcal{O}(j)$, and the tree as a whole adds $\mathcal{O}(jm)$ to the single global grid of size $n$. Provided $j\ll n$, the combined memory footprint, $\mathcal{O}(jm+n)$, is therefore orders of magnitude smaller than the $\mathcal{O}(mn)$ required by the naïve approach while achieving the same incremental information-gain calculation.
While implementing this embedded method, we observed that hash map key selection significantly impacts computation time. Using strings as keys resulted in substantial overhead due to frequent string-to-integer conversions. Even transitioning from a two-integer key to a single-integer key yielded a notable improvement in performance and is the approach we use in our implementation through a simple index-to-key conversion. Since cell value lookups occurs frequently, optimizing this efficiency significantly accelerated search tree expansion.
@@ IV-D1 Belief Map Node Embedding
By incorporating this belief map node embedding strategy, our planning framework efficiently maintains an accurate representation of the evolving belief map within the search tree. This enables the planner to greatly reduce computational overhead and create higher quality paths.
@@ IV-E Sensor Modeling
The updates to our belief map are dependent on the sensor model, which captures the performance characteristics of the perception system employed. There are many environmental factors that affect the performance of a sensor, such as visibility and range, which can be incorporated into this model [54, 55, 51].
For our approach, we formulate the sensor model as a piecewise function and is implemented as a lookup table. For a given range $r$ and binary measurement $Z$, this table maps the range of the object from the sensor to the corresponding true positive rate (TPR) $P\left(Z|X,r\right)$ and true negative rate (TNR) $P\left(\overline{Z}|\overline{X},r\right)$ of an object detection model. A typical range-based detection model will have better performance at close ranges and then tapers off as the distance increases. Using the TPR and TNR of the sensor model, we apply Bayes’ theorem to update the posterior belief of $X_{i}$ in the belief map given the measurement $Z$.
@@ IV-F Information Reward Models
The information reward function $I(\cdot)$ used in our framework is not constrained to being modular and could be time-varying or even submodular, depending on the specific requirements of the application. In this work, we have implemented an objective of reducing the overall entropy of the belief map. As defined in Section IV-D, our belief map represents the probability of the presence of an object of interest in a cell. The reward for one cell $I(X_{i})$ is calculated by finding the difference between the initial entropy and the final entropy after updating the belief with the measurements and sensor model defined in Section IV-E.
@@ IV-F2 Priority-dependent Rewards
In the case where certain areas of an environment or spatial information are more important than others, such as areas around populated regions or critical infrastructure, we introduced the ability to weight the information reward by a priority scalar, $p_{i}$, representing the priority of cell $i$. This approach provides finer control over the UAV behavior by allowing the entropy of different cells to be weighted unequally. Assigning a higher priority to one prior over another results in paths that favor observations of the higher-weighted prior. For example, this is useful when one type of object is more important than another, and the prior spatial distributions are concentrated in distinct areas of the search space (as demonstrated in Section V-C3).
@@ IV-F3 Time-dependent Rewards
In our use cases, we implement a simple decay function
@@
where $\Gamma(t)\in[\gamma,1]$, and $\gamma\in[0,1]$ is the floor value of the decay function. $\beta<0$ is the decay rate and $\gamma_{t}=(\gamma-1)/\beta$ is the time where the function output becomes the floor value $\gamma$. This gives a linear decay until the floor value.
@@ V Simulation Evaluation & Results
Figure 7: IA-TIGRIS continuously replans and refines the global path over the simulation. These snapshots show how the path adapts online to maximize information gain, with the search tree visualized to illustrate focus on high-information areas while respecting the remaining budget. For clarity, the visualized entropy map is static in the figure to highlight the evolving planned path, but in reality the belief map updates continuously as measurements are taken.
@@ V-A Simulation Framework
We use a simplified simulator, shown in Fig. 6, to enable rapid development and validation of our approach. This simulator models scenarios where a UAV is tasked with mapping the location of objects of interest in the belief map by planning paths that maximize the reduction of the entropy. The prior belief map could be informed by previous flight data, user input, or data from other sensor modalities. The visualized color for each grid cell represents the entropy of the belief that an object lies within that space, with pink regions corresponding to areas of higher entropy and blue regions indicating low entropy. The simulation employs a constant-velocity motion model for the fixed-wing UAV. To account for the effects of banking on observations, we compute the orientation change between consecutive waypoints. If this change exceeds a threshold, the UAV is considered to be banking, and the belief map is not updated during this period. This reflects the fact that the camera is primarily pointed toward the sky during banking and is also affected by increased motion blur.
@@
The UAV commanded speed is $25$ m/s, with a planning time of $10$ seconds and a the flight altitude of $50$ m. The turning radius is $100$ m, and the onboard camera is pitched $30$ degrees from horizontal with a $36.9$-degree field of view. The TPR and TNR values are $0.9$ for ranges of 0-$200$ m and then decrease linearly to $0.5$ at $600$ m. The belief map grid cells are $30$ m, and the total budget for flight distance is $15,000$ m. The belief map is updated using Bayes’ theorem to incorporate the projected camera field of view, simulated observations, and the sensor model. To accelerate testing, the simulation runs at twice the real-time speed, with results subsequently rescaled to match real-time conditions.